# Establishing Minimum Elements for Effective Vulnerability Management in AI Software

### v1

@@ II-E Exploitability
Describes how a vulnerability can be exploited and reproduced. In this field, the following details are considered:
  • Technical Complexity: Describes the technical skills and knowledge required to exploit the vulnerability in the AI system. This might include the sophistication of the techniques that must be employed, such as crafting inputs that deceive AI models (e.g., adversarial examples) or exploiting specific weaknesses in the AI’s design.
  • Required Privilege Level: Indicates whether an attacker needs elevated privileges to exploit the vulnerability or if it can be accessed with no or minimal privileges. For AI systems, this could also include access to various properties such as model parameters, data, or the inference model.
  • Specific Actions Required: Details the steps that an attacker must take to successfully exploit the vulnerability.
@@ II-F Mitigation
This section details strategies to address vulnerabilities in AI systems. It assesses vulnerabilities based on the specific AI model, data, and domain, providing tailored strategies for patching flaws and enhancing system resilience. This may include software patches, model updates, or data processing revisions. Additionally, broader system upgrades may be recommended to strengthen defenses against future threats, ensuring enhanced security and system integrity. Each entry in the mitigation section offers a comprehensive analysis of a technique, detailing its effects, advantages, and disadvantages, as well as the vulnerabilities it addresses. For example, the Adversarial Example Detection technique[16] serves as a countermeasure against adversarial machine learning attacks. TableIV presents the details of this mitigation technique.
Description The technique employs distance-based anomaly detection to identify adversarial examples using a subset of trusted data points. It divides these data points into two groups and trains a distance-based outlier detector for each group. During training, it establishes a pair of thresholds that are used to detect outliers. This approach helps in distinguishing between legitimate and adversarial inputs by analyzing the distances from these trusted points.
@@
TABLE IV: An example of a Mitigation Technique
Description: The technique employs distance-based anomaly detection to identify adversarial examples using a subset of trusted data points. It divides these data points into two groups and trains a distance-based outlier detector for each group. During training, it establishes a pair of thresholds that are used to detect outliers. This approach helps distinguish between legitimate and adversarial inputs by analyzing distances from trusted points.
TABLE V: AI Vulnerability Details
@@
Description The malicious actor first collects data similar to the target model’s training data to train multiple shadow models, which mimic the target model’s behavior. These models are then used to classify data points as ”in” (used in training) or ”out” based on their outputs. An attack model is subsequently trained on these classifications to learn the distinction between data used and not used in training. This attack model is finally applied to the target model’s outputs to infer whether specific data points were part of its training set, thus potentially exposing sensitive information if the attack is successful.
AI Model Detail Meta Data:
@@ 2. Insufficient coarsening of the prediction vector
Impact Expose individual data entries used in a model’s training set, leading to significant privacy violations and the risk of leaking confidential or personally identifiable information.
AI-Severity Score 9.0 / 10.0 (Critical). This vulnerability exposes confidential information, has Medium Exploitability, and affects a broad range of software and applications.
@@
Exploitability The malicious actor needs to have (partial) access to the training data distribution, and should be able to query the target model for multiple times as required. The model also reveals the output vector of the prediction score.
Mitigation For this vulnerability, three mitigation strategies are advised: - Restricting the Prediction Vector to Top-K Classes: This technique involves limiting the model’s output to only the top-K probable classes for each prediction. By reducing the amount of detailed information available in the output, it becomes more challenging for an attack model to determine if a data point was part of the training set. - Coarsening the Precision of the Prediction Vector: This method reduces the granularity of the output probabilities, such as rounding or binning the values, which obscures the precision of how well a data point fits the model trained with the training data. This lack of precision hinders the attack model’s ability to make accurate membership inferences. - Differential Privacy: Adding controlled noise to the data or the model’s parameters during training, differential privacy ensures that the model’s outputs are approximately the same, regardless of the inclusion or exclusion of any individual data point. This approach not only protects against membership inference attacks but also helps comply with privacy regulations and builds user trust, especially in handling sensitive data.
@@ III AI Vulnerability Disclosure Process
The AI Vulnerability Disclosure Process should be tailored to address the unique characteristics and potential impacts of vulnerabilities within artificial intelligence systems. The process can leverage the best practices of software vulnerability disclosure processes to ensure efficient, responsible, and secure handling of vulnerabilities from discovery to resolution. This requires establishing organizational policies, assigning specific roles within the organization for handling disclosures, including a dedicated security team responsible for initial assessments, communications, and coordination with external researchers. Additionally, this requires conducting a preliminary analysis to verify the reported vulnerability. This may involve replicating the issue or reviewing the AI system’s code and behavior under controlled conditions. Within the disclosure process, the potential impact of the vulnerability on the AI system and its users needs to be evaluated. Factors such as the type of data at risk, the severity of the impact on system functionality, and the likelihood of exploitation need to be considered. Additionally, the criteria for responsible disclosure of AI vulnerabilities need to be developed.
Existing vulnerability management best practices can be adopted to support collecting, analyzing and determining the minimum elements. For instance, there will be a need for forming AI CVE Naming Authorities (CNAs), these will be organizations, vendors, researchers, Bug Bounty providers that are authorized to assign CVE IDs to vulnerabilities within their scope, facilitating standardized security vulnerability reporting and management. These entities will intake vulnerability reports and assign AI-CVE IDs and determine each of the minimum elements such as #1.AI-CVE ID, #7. Affected Software and Products and #8. CPE ID, #12. References, #13. Report Date will be determined through CNAs.
Existing sources such as AI/ML ATT&CK Framework[11, 20] can be utilizes to determine information required for the Minimum Elements such as #3. Weakness Type, #4. Root Cause, #5. Impact, and #10. Mitigation. Figure1 represents an overview of AI/ML ATT&CK Framework.
Figure 1: An overview of AI/ML ATT&CK Framework[11].
@@ IV AI Bill of Materials to Support Minimum Elements
The AI Bill of Materials (AIBoM) is essential for providing deep insights into the origins and makeup of AI systems, thereby enhancing the traceability of components and the reproducibility of results. It plays a crucial role in identifying potential vulnerabilities and ensuring that ethical standards are maintained throughout the lifecycle of AI products. While there are established resources and guidelines for creating SBOM[21, 22], there is a notable absence of a standardized framework for AIBOM, highlighting a critical gap in the management of AI components. To the best of our knowledge, currently, there is no specific and comprehensive definition of AIBOM in the literature. Therefore, in this work, we propose AIBOM, a framework designed to fill this gap by providing structured guidelines and tools for documenting and managing diverse elements of AI systems. The proposed AIBOM will be used in AIVD to systematically catalog and assess vulnerabilities across AI models, enabling a more robust defense mechanism against potential threats and enhancing the overall security posture of AI deployments.
AIBOM consists of several components [23, 24], as shown in Figure2 which are detailed as follows:
@@ IV-1 Meta Data
This component is responsible for providing a comprehensive overview of the AI model’s identity and background, ensuring clear documentation of its core characteristics.
  • Name: The name of the AI model.
  • Type: Type of the core model (type of the used machine learning or deep learning algorithm)
  • Version information
  • Dependencies: Software or other models that are prerequisites.
  • Availability: Open Source, Restricted Access, Private
  • Creator of the model
  • Certification such as official approval or compliance with industry standards.
  • Release Date
  • License that is legal permissions for use, modification, and distribution of the model.
@@ IV-2 Model
: Details of the architecture and the model consisting of:
  • Foundation Model: Details about the underlying model used as a foundation.
  • Additional Model: Information about any supplementary models integrated with the base model. to the Foundation Model
  • Software Requirements
  • Hardware Requirements
  • Dependencies: Third party libraries and packages required for this model.
@@ IV-3 Data
: Details of the Data used in this model including:
  • Data Source: Description of where and how data is sourced.
  • Collection Method: How the data is gathered.
  • Quality Assessment: Processes to ensure and evaluate data quality.
  • Governance: Policies and procedures managing data access, usage, and compliance.
  • Annotation Process: Details about labeling and annotation methods.
@@ IV-4 Consideration
  • Ethical: Adherence to ethical guidelines and societal impact assessments.
  • Environmental: Analysis of energy consumption and carbon footprint.
@@ IV-5 Usage
  • Intended Usage that is the primary application
  • Out of Scope & Malicious Usage that are scenarios that are not supported or are potentially harmful
V Motivating Examples
In this section, we present a real-world example that demonstrates how an AI vulnerability is identified, modeled, and managed within our AIVD framework. This example serves to illustrate the practical application of the AIVD in addressing a specific, documented vulnerability, highlighting the framework’s effectiveness in enhancing the reliability of AI technologies. Through this case study, we aim to showcase the capabilities of the AIVD in responding to actual threats faced by AI systems in various industries.
V-1 A Real-World AI Vulnerability
In this case, as shown in FigureV, we detail a vulnerability which malicious actors can infer the presence of individual data points in a model’s training set, which allows the malicious actor to deploy a Membership Attack, potentially violating data privacy and exposing sensitive information.
@@ VI Moving Forward
We need to discuss the gaps, for instance “severity” can be part of AIVD, but how do we measure the severity of AI incidents?
As we aim to develop trustworthy and reliable AI systems, it is crucial to establish robust mechanisms for managing AI vulnerabilities. This section examines the urgent challenges and significant gaps that must be addressed to enhance the reliability and security of AI deployments, paving the way for more resilient AI frameworks.
@@ VII Conclusions
In this paper, we introduce the AIVD framework, designed to report and evaluate vulnerabilities within AI systems. We define the Minimum Elements necessary to accurately describe an AI vulnerability within this framework. Additionally, we propose AIBOM to describe and document the components and dependencies of AI systems in detail. This paper also provides examples of capturing and characterizing AI weaknesses and vulnerabilities and highlights challenges in effectively managing AI vulnerabilities. It underscores the importance of a systematic approach for identifying and mitigating potential threats in AI deployments.
VIII Acknowledgements
We thank the reviewers for their valuable comments.

### v2

@@ II-E Exploitability
Describes how a vulnerability can be exploited or reproduced. The following factors are considered:
  • Technical Complexity: Level of skill and sophistication needed, e.g., crafting adversarial inputs or exploiting model design flaws.
  • Required Privilege Level: Whether elevated privileges are needed; may include access to model parameters, training data, or inference APIs.
  • Specific Actions Required: Steps an attacker must perform to exploit the vulnerability.
@@ II-F Mitigation
Mitigation strategies in AI systems must account for the specific model, data, and domain. These may include software patches, model updates, or data preprocessing. In some cases, broader system-level changes are needed to improve resilience and prevent future threats.
Each mitigation technique includes an analysis of its purpose, strengths, limitations, and the specific vulnerabilities it addresses. For example, the Adversarial Example Detection technique [17] is used to defend against adversarial machine learning attacks. Table III outlines the details of this method.
TABLE III: An example of a Mitigation Technique
Description The technique employs distance-based anomaly detection to identify adversarial examples using a subset of trusted data points. It divides these data points into two groups and trains a distance-based outlier detector for each group. In training, it establishes a pair of thresholds that are used to detect outliers. This approach helps in distinguishing between legitimate and adversarial inputs by analyzing the distances from these trusted points.
@@
TABLE IV: AI Vulnerability Details
@@
Description The malicious actor collects data similar to the target model’s training data to train multiple shadow models that mimic the target’s behavior. These models classify points as ”in” (used) or ”out” based on outputs. An attack model is trained on these classifications to distinguish membership, then applied to the target’s outputs to infer whether specific data points were in its training set, potentially exposing sensitive information.
AI System Details Meta data:
@@ 2. Insufficient coarsening of the prediction vector
Impact Expose individual data entries used in training, leading to privacy violations and the risk of leaking confidential information.
AI-Severity Score 9.0 / 10.0 (Critical). This vulnerability exposes confidential information, has Medium Exploitability, and affects a broad range of applications.
@@
Exploitability The malicious actor needs to have (partial) access to the training data distribution and repeated query access to the target model; the model must expose full prediction vectors.
Mitigation For this vulnerability, three mitigation strategies are advised: - Restricting the Prediction Vector to Top-K Classes: Limits outputs to the top-K most probable classes, reducing information and making it harder to infer whether a data point was in the training set. - Coarsening Prediction Precision: Lowers output granularity (e.g., via rounding or binning), obscuring how closely a data point matches the model and reducing inference accuracy. - Differential Privacy: Adds controlled noise to data or parameters during training so outputs remain nearly unchanged with or without any individual record, protecting against membership inference and ensuring privacy.
@@ III AI Vulnerability Disclosure Process
The AI Vulnerability Disclosure Process must reflect the unique nature and impact of AI vulnerabilities. While it can build on existing software disclosure practices, it should address AI-specific complexity through clear policies, defined roles (e.g., a security team for triage), and verification via replication or controlled testing. The process should assess impact based on data sensitivity, functionality, and exploitability, and define clear criteria for responsible disclosure.
Existing vulnerability management practices can be adapted to support collecting, analyzing, and defining the minimum elements. For example, establishing AI CVE Naming Authorities (CNAs)—organizations, vendors, researchers, or bug bounty providers authorized to assign CVE IDs—will help standardize AI vulnerability reporting and management. These entities will intake vulnerability reports and assign AI-CVE IDs and determine each of the minimum elements such as #1.AI-CVE ID, #7. Affected Software and Products and #8. CPE ID, #12. References, #13. Report Date will be determined through CNAs.
Existing sources such as AI/ML ATT&CK Framework[14, 22] can be utilizes to determine information required for the Minimum Elements such as #3. Weakness Type, #4. Root Cause, #5. Impact, and #10. Mitigation. The framework categorizes AI security information into three primary components: Weaknesses, which describe root causes, types; Attacks, characterized by their tactics, motivation, access level, and affected models; and Mitigation, outlining defense types, their advantages, limitations, and overall effectiveness.
@@ IV AI Bill of Materials to Support Minimum Elements
AIBOM provides deep insights into the origins and composition of AI systems, enhancing component traceability and result reproducibility. It also plays a crucial role in identifying vulnerabilities and ensuring ethical standards throughout the AI product lifecycle. While resources and guidelines exist for creating SBOMs [23, 24], no standardized framework yet exists for AIBOM, revealing a critical gap in AI component management. To our knowledge, there is no comprehensive definition of AIBOM in the literature. In this work, we propose AIBOM, a framework that fills this gap by offering structured guidelines to document and manage elements of AI systems. AIBOM supports AIVD in systematically cataloging and assessing vulnerabilities across AI models, strengthening defenses and improving the reliability of AI deployments.
As shown in Figure 1, AIBOM includes several components [25, 26, 27], described below:
@@ IV-1 Meta Data
Describes the model’s identity, origin, and authorization.
  • AIBOM Generation Tool: Tool used to create the AIBOM record.
  • Creator: Individual or organization that developed the model.
  • Certification: Official approvals or compliance credentials.
  • Release Date: Date the model was published or deployed.
  • License: Legal terms governing use, modification, and distribution.
@@ IV-2 Model
Details of the model architecture and components include:
  • Source: Origin or provider of the model.
  • Availability: Public or restricted access status.
  • Foundation Model: Base model architecture used for adaptation.
  • Additional Model: Supplementary models integrated with the foundation model.
  • Weights: Pretrained or fine-tuned parameters of the model.
  • Scripts: Code used to run the model.
  • Hyperparameters: Key tuning parameters defining model behavior.
  • Configurations: Settings required for execution.
  • Domain: Application area such as healthcare, finance, or security.
  • Training Process: Procedure followed to train the model.
  • Software Requirements: Necessary frameworks or software versions.
  • Hardware Requirements: Hardware resources needed for operation.
  • Evaluation Process: Methods used to assess performance.
  • Dependencies: Libraries and packages required by the model.
@@ IV-3 Data
Details of the Data used in this model including:
  • Data Source: Origin of the dataset.
  • Availability: Either public or private
  • Collection Method: Process used for data gathering.
  • Preprocessing: Steps applied for data preparations.
  • Input/Output Format: Structure and format of data.
  • Quantitative Measures: Numerical properties of data such as size.
  • Qualitative Measures: Descriptive aspects of data such as context relevance and annotation clarity.
  • Governance: Policies managing data access and usage.
  • Annotation: Procedures for data labeling.
@@ IV-4 Consideration
Covers ethical, environmental, and operational factors influencing the responsible use of the system.
  • Ethical: Compliance with ethical standards and societal values.
  • Environmental: Broader ecological impact, e.g., data center sustainability.
  • Energy Usage: Power consumed during training and inference.
  • Carbon Footprint: CO2 emissions from energy use.
  • Risk: Potential ethical or technical vulnerabilities.
  • Mitigation: Actions taken to reduce identified risks.
  • Recommendation: Suggested improvements or best practices.
@@ IV-5 Usage
Defines the intended scope and boundaries of how the model should or should not be applied.
  • Intended Usage: Primary application.
  • Out of Scope: Scenarios beyond the defined use cases.
  • Malicious Usage: Potential misuse or harmful applications of the system.
V Motivating Example
We present a real-world case (tested on AWS and Google Cloud) illustrating how an AI vulnerability is identified, modeled, and managed within the AIVD framework. This example demonstrates AIVD’s practical use in addressing a documented vulnerability and its effectiveness in improving AI system reliability.
A Real-World AI Vulnerability: Table IV presents a real case of a membership inference attack, where a malicious actor queries an AI model to determine whether a data point was part of its training set. By exploiting prediction outputs, the attacker can infer sensitive information, violating privacy and regulatory standards. This case shows how traditional databases overlook AI-specific threats, while AIVD captures them through detailed fields like model information, severity, and mitigation, highlighting the need for a dedicated AI vulnerability database.
@@ VI Moving Forward
We must address existing gaps—for example, while “severity” can be part of AIVD, how should it be measured for AI incidents?
Building trustworthy AI systems requires robust mechanisms for managing vulnerabilities. This section outlines the main challenges and gaps to strengthen the reliability, security, and resilience of AI deployments.
@@ VII Conclusions
This paper introduces AIVD, a unified framework for managing vulnerabilities in AI systems. We define MEs for structured reporting, propose AI-CWE for AI-specific weaknesses, and present AIBOM to document AI system components. We also highlight key challenges unique to AI and illustrate how our framework enables more effective identification and mitigation of AI vulnerabilities.